OpenStudy (anonymous):

What are the solutions to the equation x2 + 4 = 148? A. –11 and 11 B. –2 and 2 C. –12 and 12 D. –4 and 4

OpenStudy (anonymous):

@Chiko_1278

jaynator495 (jaynator495):

by x2 do you mean \[\sf \Huge x^{2} \]

OpenStudy (anonymous):

i helped you already.....with this one...O.O

OpenStudy (anonymous):

First step is to rearrange the equation, getting everything on one side x^2 - 144 = 0 At this point it's pretty clear what to do Difference of Perfect Squares (x+12)(x-12) = 0 (x+12) = 0, (x-12) = 0 x = -12, 12

OpenStudy (anonymous):

thats what i told you @posingocean

jaynator495 (jaynator495):

Raise x to the 2nd power (x^2+4)-148=0 Simplify x^2+4-148 (x+12)*(x-12)=0 Solve (x+12)*(x-12)=0 x+12 = 0 Subtract 12 from both sides of the equation : x = -12 x-12 = 0 Add 12 to both sides of the equation : x = 12 12 or -12

OpenStudy (anonymous):

ik i accidentally put this on here i had a diff question